OpenStudy (anonymous):

How to find the perpendicular bisector of segment AB if A (-2,4) and B (4,0)

OpenStudy (jdoe0001):

a perpendicular bisector to that segment AB, will be another segment that makes say 90 degrees so 

OpenStudy (anonymous):

bisector -> find the midpoint. M=(1,2) is one point on the perpendicular bisector. perpendicular -> slope of AB=-2/3; slope of perpendicular line=3/2 y=mx+b y=3/2 (x-1)+2